On February 1, 2021 Marianne Co. purchased the right to remove gravel from an old rock quarry. The gravel is to be sold as roadbed for highway construction. The cost of the quarry rights was ₱164,000, with estimated salable rocks of 20,000 tons. During 2021, Marianne loaded and sold 4,000 tons of rock and estimated that 16,000 tons remained at December 31, 2021. At January 1,2022, Marianne estimated that 20,000 tons still remained. During 2022, Marianne loaded and sold 8,000 tons.

Questions:

Based on the above data, answer the following:

1. How much is the depletion in 2021?



1
Expert's answer
2021-06-17T13:20:29-0400

Solution:

a.). Depletion in 2021:


Depletion = "\frac{Total \;costs - Salvage \;value}{Estimated \;no.\; of\; units}\times No.\; of \;units\;extracted"


Total costs = 164,000

Estimated no. of units = 20,000 units

No. of units extracted in 2021 = 4,000 units


Depletion in 2021 = "\frac{164,000 - 0}{20,000}\times 4,000 = 8.2\times 4,000 = 32,800"


Depletion in 2021 = ₱32,800